OpenStudy (anonymous):

Determine the exact value of the cosecant of the quadrant angle pi/2.

OpenStudy (anonymous):

cosecant or csc is just 1 / sin \[csc(x) = \frac{1}{sin(x)}\] \[sin \left(\frac{\pi}{2}\right) = 1 \] as seen here: http://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/300px-Unit_circle_angles_color.svg.png So \[csc\left(\frac{\pi}{2}\right) = \frac{1}{1} = 1\]

OpenStudy (amorfide):

just remember 1/sin(x) = cosec 1/cos(x)= sec 1/tan=cot cos/sin=cot

OpenStudy (anonymous):

y'all rock. Just to make sure I get it...in the last step...why does pi over 2 =1/1..